OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

first they should have same slope i.e. 7

OpenStudy (luigi0210):

well if it's parallel then use the same slope, which in this case is 7 Just put it into the point-slope fomula: y-y1=m(x-x1) The x1 and y1 come from the listed points

OpenStudy (luigi0210):

then solve it, and leave it in y=mx+b

OpenStudy (anonymous):

then substitute the values in the equation (y- y1)=m(x-x1)

OpenStudy (anonymous):

(y- y1)=m(x-x1)